# Moner: Motion Correction in Undersampled Radial MRI with Unsupervised Neural Representation

### v1

@@ Abstract
Motion correction (MoCo) in radial MRI is a challenging problem due to the unpredictability of subject’s motion. Current state-of-the-art (SOTA) MoCo algorithms often use extensive high-quality MR images to pre-train neural networks, obtaining excellent reconstructions. However, the need for large-scale datasets significantly increases costs and limits model generalization. In this work, we propose Moner, an unsupervised MoCo method that jointly solves artifact-free MR images and accurate motion from undersampled, rigid motion-corrupted k-space data, without requiring training data. Our core idea is to leverage the continuous prior of implicit neural representation (INR) to constrain this ill-posed inverse problem, enabling ideal solutions. Specifically, we incorporate a quasi-static motion model into the INR, granting its ability to correct subject’s motion. To stabilize model optimization, we reformulate radial MRI as a back-projection problem using the Fourier-slice theorem. Additionally, we propose a novel coarse-to-fine hash encoding strategy, significantly enhancing MoCo accuracy. Experiments on multiple MRI datasets show our Moner achieves performance comparable to SOTA MoCo techniques on in-domain data, while demonstrating significant improvements on out-of-domain data.
@@ 1 Introduction
Recent works (Singh et al., 2024; Levac et al., 2023; 2024) propose integrating pre-trained neural networks with model-based optimization. Data priors from the pre-trained networks effectively narrow the solution space, while model-based optimization ensures reliable data consistency using MRI physical models, enabling the recovery of high-quality MR images with high data fidelity. However, pre-training the networks necessitates numerous diagnosis-quality MR images, significantly increasing reconstruction costs and often suffering from out-of-domain (OOD) problem. These limitations reduce their practicality and reliability in real-world scenarios.
@@ 2.1 Radial MRI k-space Data versus Projection Data
where $\mathcal{T}^{-1}_{\omega}\{\cdot\}$ denotes the 1D IFT operator over the variable $\omega$ and $\delta(\cdot)$ is the Dirac delta function.
@@ 2.3 Implicit Neural Representation for MRI Reconstruction
Implicit neural representation (INR) has emerged as a universal unsupervised DL framework for visual inverse problems, such as novel view synthesis (Mildenhall et al., 2021) and surface reconstruction (Wang et al., 2021). Its core concept is to leverage the learning bias inherent in neural networks (Rahaman et al., 2019; Xu et al., 2019) to regularize the underdetermined nature of inverse problems, while incorporating differentiable forward models to simulate physical processes for high data fidelity. Recently, many INR-based MRI reconstruction methods (Feng et al., 2023; Shen et al., 2022; Xu et al., 2023; Kunz et al., 2023; Feng et al., 2022; Spieker et al., 2023b; Wu et al., 2021; Huang et al., 2023; Chen et al., 2023c) have been proposed. By using Fourier transforms to simulate MRI acquisition, they can reconstruct high-quality MR images with high data fidelity. However, these INR-based methods fail to address motion-corrupted radial MRI, since they lack motion modeling and overlook the high dynamic range problem associated with k-space data. Instead, our Moner incorporates a motion model into the INR framework and introduces a new formulation for the radial MRI, fundamentally addressing the MRI MoCo problem.
@@ 4.1 Experimental Setups
Datasets and Prepossessing
The fastMRI (Knoll et al., 2020) and MoDL (Aggarwal et al., 2018) datasets are used in our experiments. For the fastMRI dataset, we extract 1,925 brain MR images with a resolution of 320$\times$320 and split them into three parts, including 1,800 slices for training set, 100 for validation set, and 25 for test set. The training and validation sets are used solely for training supervised baselines, while our Moner does not access them. For the MoDL dataset, we use 20 brain MR scans with a resolution of 256$\times$256 for additional evaluation. The acceleration factors (AF) of undersampled MRI are set to 2×2$\times$2 × and 4×4$\times$4 ×. We simulate four levels of rigid motion $\beta=\{2,5,10,15\}$ with random rotations of $[-\beta,\beta]^{\circ}$ and shifts of $[-\beta,\beta]$ mm along the X-axis and Y-axis. The MRI reconstructions under different settings (i.e., different AFs and motion ranges $\beta$) are considered as different tasks. Thus, all training and test processes are conducted independently.
@@
Four representative MRI MoCo approaches are compared, including: 1) one analytical algorithm (NuIFFT (Fessler, 2007)), 2) one iterative optimization algorithm (TV (Rudin et al., 1992)), 3) one image-based supervised DL models (DRN-DCMB (Liu et al., 2020)), and 4) one self-supervised diffusion-based DL method (Score-MoCo (Levac et al., 2023; 2024)). The Score-MoCo currently is the SOTA model for the MRI MoCo. For the reconstructed MR images, we use peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) as quantitative evaluation metrics. For the estimated motion parameters, we compute the standard deviation of absolute errors between estimations and trues, denoted by $\sigma_{\vartheta}$ and $\sigma_{\bm{\tau}}$. More details of the baselines and metrics can be found in Appendix A.1 and A.2.
Implementation Details
For the hash encoding (Müller et al., 2022) used in our model, we set its hyper-parameters as follows: base resolution $N_{\text{min}}$ = 2, maximal hash table size $T$ = 218, and resolution growth rate $b$ = 2. Our coarse-to-fine strategy sets its resolution $L$ from 4 to 16 as optimization progresses. At each iteration, we randomly sample 80 rays (i.e., $|\mathcal{R}|$ = 80 in Eq. 7). We use the Adam algorithm with default hyper-parameters (Kingma & Ba, 2014) to optimize the model. The learning rate is initialized to 0.001 and decays by half every 1,000 epochs. The total number of epochs is 4,000. It is important to note that here the hyper-parameters are determined based on 5 samples from the training set of the fastMRI dataset (Knoll et al., 2020) and are kept consistent across all other cases.
@@
fastMRI 2$\times$ $\pm$2 0.038///0.290 0.038///0.293 0.009///0.057
$\pm$5 0.038///0.309 0.038///0.313 0.010///0.131
$\pm$10 0.037///0.407 0.040///0.470 0.009///0.305
$\pm$15 0.036///0.621 0.044///0.648 0.010///0.602
4$\times$ $\pm$2 1.514///0.676 0.052///0.312 0.034///0.060
$\pm$5 1.369///0.670 0.062///0.296 0.047///0.119
$\pm$10 1.699///0.922 0.047///0.411 0.042///0.279
$\pm$15 2.513///1.520 0.050///0.532 0.043///0.812
MoDL 2$\times$ $\pm$2 0.037///0.275 0.389///0.409 0.009///0.058
$\pm$5 0.040///0.276 0.838///0.720 0.009///0.144
$\pm$10 0.042///0.371 0.748///1.263 0.008///0.286
$\pm$15 0.040///0.553 0.096///1.448 0.009///0.570
4$\times$ $\pm$2 0.167///0.348 1.946///0.824 0.019///0.065
$\pm$5 0.276///0.377 1.125///0.782 0.021///0.163
$\pm$10 0.257///0.494 1.642///1.242 0.022///0.324
$\pm$15 0.506///0.824 1.721///1.797 0.019///0.578
Table 1: Quantitative results ($\sigma_{\vartheta}/\sigma_{\bm{\tau}}$) of motion parameters by compared methods on the fastMRI and MoDL datasets. Here the “AF” and “MR” represent acceleration rate and motion range, respectively. The best and second performances are highlighted in blod and underline, respectively.
@@ 4.2 Main Results
Dataset AF MR Analy. Optim. Sup. Self-Sup. Unsup.
@@
fastMRI 2$\times$ $\pm$2 27.16///0.537 29.53///0.860 29.80///0.836 33.94///0.866 32.64///0.915
$\pm$5 22.93///0.364 29.53///0.859 25.63///0.716 33.64///0.849 32.62///0.914
$\pm$10 20.46///0.274 29.46///0.859 22.62///0.624 33.52///0.851 32.49///0.915
$\pm$15 19.55///0.251 29.44///0.859 21.07///0.532 33.69///0.857 32.50///0.913
4$\times$ $\pm$2 25.99///0.460 25.12///0.613 29.18///0.812 32.33///0.813 31.49///0.888
$\pm$5 22.30///0.308 25.56///0.645 25.18///0.715 32.03///0.800 31.07///0.879
$\pm$10 19.81///0.221 25.41///0.639 22.01///0.570 32.34///0.809 31.20///0.882
$\pm$15 19.20///0.206 25.12///0.610 21.22///0.535 32.22///0.804 31.07///0.881
MoDL 2$\times$ $\pm$2 28.51///0.513 31.33///0.888 29.66///0.793 32.76///0.780 34.56///0.952
$\pm$5 25.15///0.336 31.32///0.887 25.79///0.613 32.55///0.760 34.54///0.952
$\pm$10 23.31///0.256 31.29///0.887 23.98///0.549 32.55///0.770 34.57///0.952
$\pm$15 22.89///0.246 31.26///0.886 22.52///0.356 32.49///0.762 34.32///0.949
4$\times$ $\pm$2 27.59///0.420 30.69///0.861 28.78///0.696 28.36///0.581 33.50///0.938
$\pm$5 24.39///0.257 30.30///0.830 25.67///0.578 29.08///0.608 33.24///0.933
$\pm$10 22.91///0.197 30.26///0.835 23.98///0.382 29.70///0.650 33.21///0.932
$\pm$15 22.40///0.185 29.81///0.811 22.72///0.371 28.51///0.608 33.25///0.933
Table 2: Quantitative results (PSNR/SSIM) of MR images by compared methods on the fastMRI and MoDL datasets. Here the “AF” and “MR” represent acceleration rate and motion range, respectively. The best and second performances are highlighted in blod and underline, respectively.
@@
Table 2 presents the quantitative comparisons of MR images reconstructed by our Moner and the baselines. On the fastMRI dataset, both Score-MoCo and our Moner produce comparable and satisfactory results, significantly outperforming the other baselines. However, when applied to the unseen MoDL dataset, the Score-MoCo model suffers from severe performance degradation due to the OOD problem, even falling behind the TV algorithm at AF = 4$\times$. This performance trend in MR image reconstruction mirrors the results in MoCo accuracy. In contrast, our Moner consistently achieves robust and high-quality reconstructions across both datasets. Fig. 4 shows the qualitative results. The analytical NuIFFT produces severe artifacts. The image-based supervised DRN-DCMB yields smooth but fuzzy MRI results. The TV algorithm, relying on local image smoothness, reduces motion artifacts but introduces strong cartoon-like features and loses many image details. While Score-MoCo recovers high-quality MR images on the fastMRI dataset, its reconstructions on the unseen MoDL dataset contain many artifacts due to inaccurate MoCo cuased by the OOD problem. Visually, our Moner consistently achieves high-quality MRI reconstructions across both datasets, further confirming its robustness and superiority over SOTA MoCo methods. Fig. 3 compares the reconstruction speeds of our method with two iterative methods (TV and Score-MoCo) using a single NVIDIA RTX 4070 Ti GPU on the fastMRI dataset. Our Moner achieves the fastest reconstruction speed, being more than 5$\times$ and 6$\times$ faster than the SOTA Score-MoCo model. Additional visual results are provided in Appendix A.3.
Figure 4: Qualitative results of MR images by compared methods on two test samples (##\##10 and ##\##19) of the fastMRI and MoDL datasets for AF = 2$\times$ and MR = $\pm$5.

### v2

@@ Abstract
Motion correction (MoCo) in radial MRI is a particularly challenging problem due to the unpredictability of subject movement. Current state-of-the-art (SOTA) MoCo algorithms often rely on extensive high-quality MR images to pre-train neural networks, which constrains the solution space and leads to outstanding image reconstruction results. However, the need for large-scale datasets significantly increases costs and limits model generalization. In this work, we propose Moner, an unsupervised MoCo method that jointly reconstructs artifact-free MR images and estimates accurate motion from undersampled, rigid motion-corrupted k-space data, without requiring any training data. Our core idea is to leverage the continuous prior of implicit neural representation (INR) to constrain this ill-posed inverse problem, facilitating optimal solutions. Specifically, we integrate a quasi-static motion model into the INR, granting its ability to correct subject’s motion. To stabilize model optimization, we reformulate radial MRI reconstruction as a back-projection problem using the Fourier-slice theorem. Additionally, we propose a novel coarse-to-fine hash encoding strategy, significantly enhancing MoCo accuracy. Experiments on multiple MRI datasets show our Moner achieves performance comparable to SOTA MoCo techniques on in-domain data, while demonstrating significant improvements on out-of-domain data.
@@ 1 Introduction
Recent works (Singh et al., 2024; Levac et al., 2023; 2024) propose integrating pre-trained neural networks with model-based optimization. Data priors from the pre-trained networks effectively constrain the solution space, while model-based optimization ensures reliable data consistency using MRI physical models, enabling the recovery of high-quality MR images with high data fidelity. However, pre-training the networks necessitates numerous diagnosis-quality MR images, significantly increasing reconstruction costs. Although these models demonstrate greater robustness compared to supervised methods, they remain vulnerable to out-of-domain (OOD) issues. These limitations undermine their practicality and reliability in real-world applications.
@@ 2.1 Radial MRI k-space Data versus Projection Data
where $\mathcal{T}^{-1}_{\omega}\{\cdot\}$ denotes the 1D IFT operator over the variable $\omega$ and $\delta(\cdot)$ is the Dirac delta function. Eq. 2 implies that the radial MRI solving (i.e., ${\bm{k}}\rightarrow{\bm{f}}$) can be theoretically reformulated as a back-projection problem (i.e., ${\bm{g}}\rightarrow{\bm{f}}$) using the 1D IFT operator. A recent study on cardiac MRI acceleration (Catalán et al., 2023) introduces the Fourier-slice theorem to optimize a neural field directly from raw k-space data, showing promising potential. Similarly, but with a distinct focus, our work uses the Fourier-slice theorem to reformulate radial MRI as a back-projection problem, fundamentally addressing the high-dynamic range issue and stabilizing optimization.
@@ 2.3 Implicit Neural Representation for MRI Reconstruction
Implicit neural representation (INR) has emerged as a universal unsupervised DL framework for visual inverse problems, such as novel view synthesis (Mildenhall et al., 2021) and surface reconstruction (Wang et al., 2021). Its core concept is to leverage the learning bias inherent in neural networks (Rahaman et al., 2019; Xu et al., 2019) to regularize the underdetermined nature of inverse problems, while incorporating differentiable forward models to simulate physical processes for high data fidelity. Recently, many INR-based MRI reconstruction methods (Feng et al., 2023; Shen et al., 2022; Xu et al., 2023; Kunz et al., 2023; Feng et al., 2022; Spieker et al., 2023b; Wu et al., 2021; Huang et al., 2023; Chen et al., 2023c; Catalán et al., 2023) have been proposed. By using Fourier transforms to simulate MRI acquisition, they can reconstruct high-quality MR images with high data fidelity. However, these INR-based methods fail to address motion-corrupted radial MRI, since they lack motion modeling. Moreover, these methods either entirely overlook the high dynamic range problem in k-space data or specifically design a relative $\ell_{2}$ loss to alleviate it (Huang et al., 2023; Spieker et al., 2023b; Feng et al., 2022). Instead, our Moner incorporates a motion model into the INR framework and introduces a new formulation for the radial MRI, fundamentally addressing the MRI MoCo problem.
@@ 4.1 Experimental Setups
Datasets
The fastMRI (Knoll et al., 2020) and MoDL (Aggarwal et al., 2018) datasets are used in our experiments. For the fastMRI dataset, we first extract 1,925 brain MR slices with image sizes of 320$\times$320 and an image spacing of 1$\times$1 mm2 from different subjects along the axial direction. These slices include three contrasts of T1w, T2w, and FLAIR. We then split them into three parts, including 1,800 slices for training set, 100 for validation set, and 25 for test set. The training and validation sets are used solely for training supervised baselines, while our Moner does not access them. For the MoDL dataset, we use 20 T2w brain MR slices with image sizes of 256$\times$256 and an image spacing of 1$\times$1 mm2 from 20 subjects along the sagittal direction for additional test.
Pre-processing
We use a 2D radial sampling pattern with the golden-angle acquisition scheme. Each spoke has a length of 511, corresponding to an imaging FOV of 511$\times$511 mm2. The fully sampled radial k-space data thus consists of approximately 720 views. For undersampled MRI, we set acceleration factors (AFs) to 2×2$\times$2 × and 4×4$\times$4 ×, corresponding to 360 and 180 views of undersampled radial k-space data, respectively. Following the motion simulation pipeline used in previous studies (Levac et al., 2023; 2024; Spieker et al., 2023a), we first divide all spokes into 18 motion stages, where spokes within the same motion stage share the same motion trajectory. Then, we simulate four levels of rigid motion $\beta=\{2,5,10,15\}$ with random rotations of $[-\beta,\beta]^{\circ}$ and shifts of $[-\beta,\beta]$ mm along the X-axis and Y-axis. The MRI reconstructions under different settings (i.e., different AFs and motion ranges $\beta$) are considered as different tasks. Thus, all training and test processes are conducted independently.
@@
Four representative MRI MoCo approaches are compared, including: 1) one analytical algorithm (NuIFFT (Fessler, 2007)), 2) one iterative optimization algorithm (TV (Rudin et al., 1992)), 3) one image-based supervised DL models (DRN-DCMB (Liu et al., 2020)), and 4) one self-supervised diffusion-based DL method (Score-MoCo (Levac et al., 2023; 2024)). The Score-MoCo currently is the SOTA model for the MRI MoCo. For the reconstructed MR images, we use peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) as quantitative evaluation metrics. For the estimated motion parameters, we compute the standard deviation of absolute errors between estimations and trues, denoted by $\sigma_{\vartheta}$ and $\sigma_{\bm{\tau}}$. More details of the baselines and metrics can be found in Appendix A.2 and A.3.
Table 1: Quantitative results (Mean in $\sigma_{\vartheta}/\sigma_{\bm{\tau}}$) of motion parameters by compared methods on the fastMRI and MoDL datasets. Results of t-test statistical tests comparing our Moner to baselines are denoted by ∗∗ ($p$-value $<$ 0.01), ($p$-value $<$ 0.05), and (not significant, $p$-value $\geq$ 0.05). Here the “AF” and “MR” represent acceleration rate and motion range, respectively. The best performances are highlighted in bold.
@@
fastMRI 2$\times$ $\pm$2 0.038∗∗///0.290∗∗ 0.038∗∗///0.293∗∗ 0.009///0.057
$\pm$5 0.038∗∗///0.309∗∗ 0.038∗∗///0.313∗∗ 0.010///0.131
$\pm$10 0.037∗∗///0.407 0.040∗∗///0.470 0.009///0.305
$\pm$15 0.036∗∗///0.621 0.044∗∗///0.648 0.010///0.602
4$\times$ $\pm$2 1.514∗∗///0.676∗∗ 0.052///0.312∗∗ 0.034///0.060
$\pm$5 1.369∗∗///0.670∗∗ 0.062///0.296∗∗ 0.047///0.119
$\pm$10 1.699∗∗///0.922∗∗ 0.047///0.411∗∗ 0.042///0.279
$\pm$15 2.513∗∗///1.520∗∗ 0.050///0.532 0.043///0.812
MoDL 2$\times$ $\pm$2 0.037∗∗///0.275∗∗ 0.389///0.409∗∗ 0.009///0.058
$\pm$5 0.040∗∗///0.276∗∗ 0.838///0.720∗∗ 0.009///0.144
$\pm$10 0.042∗∗///0.371 0.748///1.263∗∗ 0.008///0.286
$\pm$15 0.040∗∗///0.553 0.096∗∗///1.448∗∗ 0.009///0.570
4$\times$ $\pm$2 0.167///0.348∗∗ 1.946∗∗///0.824∗∗ 0.019///0.065
$\pm$5 0.276///0.377∗∗ 1.125∗∗///0.782∗∗ 0.021///0.163
$\pm$10 0.257///0.494 1.642///1.242∗∗ 0.022///0.324
$\pm$15 0.506///0.824 1.721///1.797∗∗ 0.019///0.578
Implementation Details
For our Moner, we employ the hash encoding (Müller et al., 2022) followed by two fully-connected (FC) layers with a width of 128 to implement the MLP network $\mathcal{M}_{\Phi}$. The first FC layer is followed by a ReLU activation, while the second one (i.e., , the output layer) has no activation. For the hash encoding (Müller et al., 2022) used in our model, we set its hyper-parameters as follows: base resolution $N_{\text{min}}$ = 2, maximal hash table size $T$ = 218, and resolution growth rate $b$ = 2. Our coarse-to-fine strategy sets its resolution $L$ from 4 to 16 as optimization progresses. At each iteration, we randomly sample 80 rays (i.e., $|\mathcal{R}|$ = 80 in Eq. 7). We use the Adam algorithm with default hyper-parameters (Kingma & Ba, 2014) to optimize the model. The learning rate is initialized to 0.001 and decays by half every 1,000 epochs. The total number of epochs is 4,000. Note that here the hyper-parameters are determined based on 5 samples from the training set of the fastMRI dataset (Knoll et al., 2020) and are kept consistent across all other cases.
@@ 4.2 Main Results
Table 2: Quantitative results (Mean±$\pm$±STD in PSNR) of MR images by compared methods on the fastMRI and MoDL datasets. Results of t-test statistical tests comparing our Moner to baselines are denoted by ∗∗ ($p$-value $<$ 0.01), ($p$-value $<$ 0.05), and (not significant, $p$-value $\geq$ 0.05). Here the “AF” and “MR” represent acceleration rate and motion range, respectively. The best and second performances are highlighted in bold and underline, respectively.
Dataset AF MR Analy. Optim. Sup. Unsup.
@@
fastMRI 2$\times$ $\pm$2 27.16$\pm$2.33∗∗ 29.53$\pm$2.24∗∗ 29.80$\pm$2.74∗∗ 33.94$\pm$2.29 32.64$\pm$2.65
$\pm$5 22.93$\pm$2.06∗∗ 29.53$\pm$2.19∗∗ 25.63$\pm$2.57∗∗ 33.64$\pm$2.22 32.62$\pm$2.61
$\pm$10 20.46$\pm$2.02∗∗ 29.46$\pm$2.27∗∗ 22.62$\pm$2.40∗∗ 33.52$\pm$2.55 32.49$\pm$2.78
$\pm$15 19.55$\pm$2.07∗∗ 29.44$\pm$2.28∗∗ 21.07$\pm$2.13∗∗ 33.69$\pm$2.38 32.50$\pm$2.65
4$\times$ $\pm$2 25.99$\pm$2.01∗∗ 25.12$\pm$3.91∗∗ 29.18$\pm$2.44∗∗ 32.33$\pm$2.39 31.49$\pm$2.51
$\pm$5 22.30$\pm$2.15∗∗ 25.56$\pm$3.72∗∗ 25.18$\pm$2.53∗∗ 32.03$\pm$2.36 31.07$\pm$2.58
$\pm$10 19.81$\pm$1.96∗∗ 25.41$\pm$4.05∗∗ 22.01$\pm$2.36∗∗ 32.34$\pm$2.25 31.20$\pm$2.58
$\pm$15 19.20$\pm$2.03∗∗ 25.12$\pm$3.98∗∗ 21.22$\pm$2.23∗∗ 32.22$\pm$2.38 31.07$\pm$2.74
MoDL 2$\times$ $\pm$2 28.51$\pm$1.19∗∗ 31.33$\pm$1.11∗∗ 29.66$\pm$1.04∗∗ 33.34$\pm$3.54 34.56$\pm$0.92
$\pm$5 25.15$\pm$0.96∗∗ 31.32$\pm$1.11∗∗ 25.79$\pm$0.73∗∗ 32.45$\pm$4.27 34.54$\pm$0.90
$\pm$10 23.31$\pm$1.05∗∗ 31.29$\pm$1.11∗∗ 23.98$\pm$1.10∗∗ 32.30$\pm$3.53 34.57$\pm$0.96
$\pm$15 22.89$\pm$0.97∗∗ 31.26$\pm$1.07∗∗ 22.52$\pm$1.04∗∗ 33.90$\pm$1.28 34.32$\pm$0.91
4$\times$ $\pm$2 27.59$\pm$1.08∗∗ 30.70$\pm$1.51∗∗ 28.78$\pm$1.12∗∗ 28.45$\pm$3.04∗∗ 33.50$\pm$1.03
$\pm$5 24.39$\pm$0.95∗∗ 30.30$\pm$2.00∗∗ 25.67$\pm$0.91∗∗ 30.09$\pm$2.55∗∗ 33.24$\pm$0.88
$\pm$10 22.91$\pm$1.01∗∗ 30.26$\pm$1.93∗∗ 23.98$\pm$0.87∗∗ 30.33$\pm$3.48∗∗ 33.21$\pm$1.06
$\pm$15 22.40$\pm$0.92∗∗ 29.81$\pm$1.87∗∗ 22.72$\pm$0.96∗∗ 29.94$\pm$3.19∗∗ 33.25$\pm$0.81
@@
Table 2 presents the quantitative comparisons (PSNR) of MR images reconstructed by our Moner and the baselines. The SSIM results are provided in Table 7. On the fastMRI dataset, both Score-MoCo and our Moner produce comparable and satisfactory results, significantly outperforming the other baselines. However, when applied to the unseen MoDL dataset, the Score-MoCo model suffers from severe performance degradation due to the OOD problem, even falling behind the TV algorithm at AF = 4$\times$. This performance trend in MR image reconstruction mirrors the results in MoCo accuracy. In contrast, our Moner consistently achieves robust and high-quality reconstructions across both datasets. Fig. 4 shows the qualitative results. The analytical NuIFFT produces severe artifacts. The image-based supervised DRN-DCMB yields smooth but fuzzy MRI results. The TV algorithm, relying on local image smoothness, reduces motion artifacts but introduces strong cartoon-like features and loses many image details. While Score-MoCo recovers high-quality MR images on the fastMRI dataset, its reconstructions on the unseen MoDL dataset contain many artifacts due to inaccurate MoCo cuased by the OOD problem. Visually, our Moner consistently achieves high-quality MRI reconstructions across both datasets, further confirming its robustness and superiority over SOTA MoCo methods. Fig. 3 compares the reconstruction speeds of our method with two iterative methods (TV and Score-MoCo) using a single NVIDIA RTX 4070 Ti GPU on the fastMRI dataset. Our Moner achieves the fastest reconstruction speed, being more than 5$\times$ and 6$\times$ faster than the SOTA Score-MoCo model. Additional visual results are provided in Appendix A.4.
Figure 4: Qualitative and quantitative results of MR images by compared methods on two test samples (##\##22 and ##\##9) of the fastMRI and MoDL datasets for AF = 2$\times$ and MR = $\pm$5.